OpenStudy (anonymous):

A motel clerk counts $1 and $10 bills at the end of the day. He finds that he has a total of 49 bills having a combined value of $148. Find the number of bills of each denomination that he has.

OpenStudy (anonymous):

Let x be the number of $1 and y be the number of $10. x + y = 49 x + 10y = 148 From the first, we have that y = 49 - x. Substitute this into the second and solve for x. THen substitute this value fo x into the first to get y.